At a pressure of 3.02 atm, a vessel takes up 23.4 liters of space. What pressure must this container be at for it to take up 54.5 liters of space?


1
Expert's answer
2021-01-28T05:13:05-0500

Assuming that the type of matter in question is gaseous; according to Boyle's Law, the pressure (p) of a given quantity of a gas varies inversely with its volume(v) at a constant temperature.


That is "p\alpha v"

"pv =k"


"p1v1 = p2v2"


"p2=\dfrac{p1v1}{v2}"


"p2=\dfrac{3.02atmx23.4L}{54.5L}=1.30atm"